OpenStudy (anonymous):

4. Determine the molecular formulas for compounds having the following empirical formula and molar mass: CH4O; experimental molar mass 128 g/mol

OpenStudy (anonymous):

please

OpenStudy (jebonna):

The 'experimental molar mass' is the molar mass you will need to be aiming for to get the molecular formula. So you need to get to a compound that will have a molar mass of 128 g/mol Firstly, find the molar mass of the compound you have already in empirical formula. C + 4H + O 12 + 4 + 16 = 32 g/mol So your empirical formula compound has 32 g/mol, but you need to find the compound that has 128 g/mol. So now, you will need to do: the molar mass you need / molar mass of empirical formula 128/32 = 4 So 32 goes into 128 FOUR times. This means you will need to times your empirical formula (CH4O) by 4, to get the compound with a molar mass of 128 g/mol. \[C \times 4 = C _{4}\] \[H _{4} \times 4 = H_{16}\] \[O \times 4 = O _{4}\] Now, put it all together in the same order: \[C _{4}H _{16}O _{4}\] That will be your molecular formula with a molar mass of 128 g/mol I hope this helps! :)